OpenStudy (anonymous):

which of the following is a polynomial with roots -sqrt5 ,sqrt5^-3? x^3-2x^2-3x+6 x^3+3x^2-5x-15 x^3-3x^2-5x+15 x^3+2x^2-3x-6

OpenStudy (anonymous):

i think its c

undeadknight26 (undeadknight26):

i don't think i did polynomial yet let me find out how to do it...

OpenStudy (anonymous):

oh ok

OpenStudy (solomonzelman):

\[-\sqrt{5}~~~~~~~~~~and~~~~~~~~\sqrt[5]{-3}?\]

undeadknight26 (undeadknight26):

oi oi oi...im sorry this is not something i can learn in 10 minutes...

OpenStudy (anonymous):

nope its \[-\sqrt{5} \sqrt{5} -3\]

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

\[\sqrt{5},~~~~-\sqrt{5}~~~~~~and~~~-3~~?\]

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

Set each equal zero, and factor, this is way to get it.

OpenStudy (anonymous):

so im wrong??

OpenStudy (solomonzelman):

I don;t know, I'll check.

OpenStudy (anonymous):

i got c. but if im wrong ill do it aggain

OpenStudy (solomonzelman):

x^3-3x^2-5x+15=0 x^2(x-3)-5(x-5)=0 (x^2-5)(x-3)=0 x^2-5=0 +-sqrt5 x-3=0 3

OpenStudy (solomonzelman):

close, everything beside -3.

OpenStudy (solomonzelman):

Option B. x^3+3x^2-5x-15=0 x^2(x+3)-5(x+3)=0 (x^2-5)(x+3)=0 B is it, see why?

OpenStudy (anonymous):

yea i do. can u help me with 2 more? there easy

OpenStudy (solomonzelman):

I can try, I'm not very good ta math...go!

OpenStudy (anonymous):

which polynomial identity will prove that 37=64-27? difference squares difference of cubes sum of cubes square of a binomial

OpenStudy (solomonzelman):

What what?! Honestly your teacher is not feeling well, or whatever it is.... Idk, I never had to prove such an obvious statement.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

is it d?

OpenStudy (solomonzelman):

I honestly don;t know I would tell you if I did.

OpenStudy (anonymous):

oh ok thnk

OpenStudy (solomonzelman):

Yeah sorry abt that.

OpenStudy (solomonzelman):

I can help with something else or at least try to.